OpenStudy (anonymous):

Let u = <5, 6>, v = <-2, -6>. Find -2u + 5v. <-20, -42> <-20, 0> <0, 18> <-6, 0>

OpenStudy (xapproachesinfinity):

-2u you just multiply the components of u by -2 same for 5 and then add component by component

OpenStudy (anonymous):

how so?

jimthompson5910 (jim_thompson5910):

u = <5, 6> -2*u = -2*<5, 6> -2*u = <-2*5, -2*6> ... multiply EVERY component by -2 -2*u = <-10, -12> use the same idea to find 5*v

OpenStudy (anonymous):

is it the first one

jimthompson5910 (jim_thompson5910):

what did you get for 5*v ?

OpenStudy (anonymous):

-10,-30

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

and you have the correct final answer

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np